there are 4 flowers and 2 vases. how many different combinations can be created using one flower and one vase.
4, 6, 8 or 16

Answer by RAY100(1637)

This is counting principal, m things taken n ways = m*n
.
2 vases taken 4 ways = 2*4 = 8 combinations
.